a value of b for which the eq.


                x^2 + bx -1=0


                x^2 +x +b =0


have one common root

Grade:

1 Answers

jitender lakhanpal
10 years ago

hi taru

  Let the two quadratic equations be  ax² + bx + c =0 and a1x² + b1x + c1 =0 .

then the condition that one of the root is common

(bc1-b1c)(ab1- a1b) = (ca1-ac1) 2

compare with the given equations  find the value of a , a1 , b , b1 , c , c1

 

We are  here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

jitender

askiitians expert